# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

### v4

@@ Abstract.
Next-generation datacenters require highly efficient network load balancing to manage the growing scale of artificial intelligence (AI) training and general datacenter traffic. However, existing Ethernet-based solutions, such as Equal Cost Multi-Path (ECMP) and oblivious packet spraying (OPS), struggle to maintain high network utilization due to both increasing traffic demands and the expanding scale of datacenter topologies, which also exacerbate network failures. To address these limitations, we propose ARCANE, a lightweight decentralized per-packet adaptive load balancing algorithm designed to optimize network utilization while ensuring rapid recovery from link failures. ARCANE adapts to network conditions by caching good-performing paths. In case of a network failure, ARCANE re-routes traffic away from it in less than 100 microseconds. ARCANE is designed to be deployed with next-generation out-of-order transports, such as Ultra Ethernet, and uses less than 25 bytes of per-connection state regardless of the topology size. We extensively evaluate ARCANE in large-scale simulations and FPGA-based NICs.
@@ 1. Introduction
Moreover, recent works have shown that link failures drastically impact both training times and economic costs. A single link failure can have 20$\approx$ 20$\times$ 20 × higher cost impact in distributed training workloads than in cloud workloads (Qian et al., 2024; Gangidi et al., 2024). This observation, along with the increasing scale these systems are growing at, highlights the need for a transport layer with a load balancing scheme that can adapt near-instantly, e.g., within a few round-trip times (RTTs), to network failures and, consequently, bandwidth asymmetries across the topology.
@@
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SDR (Shalev et al., 2020), UE (Lu et al., 2018), Falcon (goo, 2023) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing algorithm, Adaptive Routing with Caching and Aware Network Exploration aka ARCANE. ARCANE caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
ARCANE does not require any specific hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). ARCANE requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
We extensively evaluate ARCANE in simulation and augment the simulation findings with real hardware results in Section 4. We deploy ARCANE in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, ARCANE consistently outperforms state-of-the-art load balancing algorithms. ARCANE outperforms ECMP and OPS by up to 6×6$\times$6 × and 1.25×1.25$\times$1.25 × in symmetric networks, and by up to 4.5×4.5$\times$4.5 × and 2×2$\times$2 × in asymmetric networks, outperforming OPS by as much as 100×100$\times$100 × during short-term transient link failures.
2. ARCANE Building Blocks
In this section, we introduce the fundamental building blocks necessary to understand ARCANE logic. We first describe key congestion signals, such as ECN and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
@@ 2.1. Congestion Signals
In ARCANE, we use ECN as the congestion signal to detect path congestion due to its simplicity and widespread adoption (Hoefler et al., 2023; Nvidia, 2024). Since ECN is not marked for packets when the queue is smaller than $K_{min}$, ECN effectively filters out cases of minor queuing due to packet collisions across multiple hops, while identifying true congestion at a single bottleneck. In contrast, delay-based signals struggle to differentiate between these scenarios unless enhanced by advanced switch features, such as in-network telemetry (INT) (Wang et al., 2023).
@@
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). On current generation networks, ARCANE can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
@@
rely on congestion signals, such as ECN marks or packet loss, to adjust a flow’s sending rate or window with the goal of maintaining high network utilization while preventing queue build up. ARCANE is designed to work well with any CC algorithm as long as they support receiving and acknowledging packets out-of-order for a given message. In particular, we show later how ARCANE works well with EQDS, a variant of DCTCP, and a proprietary internal CC algorithm (Olteanu et al., 2022; Alizadeh et al., 2010).
@@ 2.2. Load Balancing
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). We leverage EVs in ARCANE to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
@@
3. ARCANE
ARCANE is a load balancing algorithm that relies on simple and memory-efficient endpoint logic. The key idea behind ARCANE is straightforward: when congestion is detected on a certain path, we explore alternative paths while caching and reusing paths with little to no congestion. Specifically, ARCANE uses a circular buffer of a fixed size to cache EVs of uncongested paths. By design, ARCANE can be implemented in NIC hardware or firmware with minimal memory/area footprint. ARCANE does not need any switch support besides ECMP hashing and ECN marking. Placing ARCANE at the endpoint not only simplifies deployment but also provides a broader, end-to-end view of the network, unlike fabric-side approaches which are limited to local visibility. Algorithm 1 details pseudocode for a ARCANE sender receiving an ACK and detecting a failure. Algorithm 2 describes pseudocode for a ARCANE sender sending a data packet.
@@ 3.1. Core Logic: Path Exploration and Reuse
During the first Bandwidth-Delay Product (BDP) worth of packets of a new or idle connection, a ARCANE sender explores random entropies from the EVS. This exploration is necessary because, initially, there is no fresh knowledge about the network’s state. In this warm up phase, ARCANE operates similarly to OPS.
@@
When an ACK arrives at the sender, if it is not ECN-marked, the EV it carries is cached in the circular buffer, and its validity bit is set to 1. Otherwise, if the ACK is ECN-marked, ARCANE does not cache the EV and discards it. When set, the validity bit indicates that an entropy has not been used after it has been added to the circular buffer. When sending a data packet out, ARCANE first checks if there are any valid EVs in its buffer. In case there is any valid EV, ARCANE reuses the oldest valid EV from the circular buffer and resets its validity bit. Otherwise, ARCANE explores a random EV from the EVS.
The circular buffer in ARCANE ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
When a network, running without ARCANE, experiences any kind of transient (e.g., link flap) or persistent unrecoverable failure (e.g., a link or switch failure), it will take the system some time to recover from it: ranging from several milliseconds to update the ECMP routing group to several seconds if a reboot is needed, and more if a swap is needed (Anbiah and Sivalingam, 2022; Iselt et al., 2004).
@@
By default, ARCANE Core Logic would already limit utilizing failing paths since it would only recycle paths that are still active and returning ACKs without ECN. However, if the ARCANE buffer was empty, it would select a random EV that could potentially map to a failing path. To further improve the performance of ARCANE and prevent further losses, we introduce freezing mode. ARCANE enters this mode when detecting failures via indirect feedback from the network. ARCANE uses a timeout heuristic (Section 2.1) that can be enhanced with packet trimming (Appendix A) as a natural feedback from the network to detect failures along a path.
When in freezing mode, ARCANE:
@@
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that ARCANE will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we use random EVs to allow ARCANE to explore new paths and assess whether we detect new packets failing or not. This prevents ARCANE from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if ARCANE exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), ARCANE would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, ARCANE can be conservative and can safely enter freezing mode.
Algorithm 1 ARCANE logic upon ACK receive and failure detection.
@@
9:    if not ARCANEBuffer[head].isValid then
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Algorithm 2 ARCANE logic on send datapath.
1$\triangleright$ Variables already listed in Algorithm 1
@@
[…]
7:    $\triangleright$ Must be in freezing mode.
[…]
@@
15:    if ARCANEBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) or exploreCounter then
[…]
@@
19:        $dataPacket.ev=\textit{getNextEV()}$()
@@
3.3. ARCANE Design Advantages
@@
The simplicity of ARCANE enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. ARCANE works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing.
@@
A key advantage of ARCANE is that it does not need to track per-EV metrics and statistics. As will be discussed in Section 4.5.2, achieving good performance with OPS requires a relatively large EVS. If OPS were to maintain metrics for each EV, the memory overhead would be excessive for a hardware NIC implementation, e.g., 64 Kib to store 1 bit per entropy value for an EVS with 64K EVs. However, ARCANE only needs a fixed number of bytes in memory. More specifically, as detailed in Table 1, ARCANE requires only around $\approx 25$ bytes. Moreover, even when constrained to a small EVS, ARCANE is still able to perform well (Section 4.5.2), which can further reduce ARCANE’ memory footprint by 1 byte since Table 1 assumed 16 bits per EV.
There is a subtle observation as to why ARCANE can achieve a great performance without needing a lot of state: while the ARCANE buffer is used to cache good entropies, it is really only useful in certain scenarios like when receiving a burst of ACKs or during freezing mode. In reality, most of ARCANE’ state is on the wire, stored in the inflight data and ACK packets, which will inform ARCANE about the good paths in the network.
@@
Table 1. Per-connection memory footprint of ARCANE.
@@
The general approach of ARCANE is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.

### v5

# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

@@ Abstract.
Next-generation datacenters require highly efficient network load balancing to manage the growing scale of artificial intelligence (AI) training and general datacenter traffic. However, existing Ethernet-based solutions, such as Equal Cost Multi-Path (ECMP) and oblivious packet spraying (OPS), struggle to maintain high network utilization due to both increasing traffic demands and the expanding scale of datacenter topologies, which also exacerbate network failures. To address these limitations, we propose REPS, a lightweight decentralized per-packet adaptive load balancing algorithm designed to optimize network utilization while ensuring rapid recovery from link failures. REPS adapts to network conditions by caching good-performing paths. In case of a network failure, REPS re-routes traffic away from it in less than 100 microseconds. REPS is designed to be deployed with next-generation out-of-order transports, such as Ultra Ethernet, and uses less than 25 bytes of per-connection state regardless of the topology size. We extensively evaluate REPS in large-scale simulations and FPGA-based NICs.
@@ 1. Introduction
Moreover, recent works have shown that link failures drastically impact both training times and economic costs. A single link failure can have $\approx 20\times$ higher cost impact in distributed training workloads than in cloud workloads (Qian et al., 2024; Gangidi et al., 2024). This observation, along with the increasing scale these systems are growing at, highlights the need for a transport layer with a load balancing scheme that can adapt near-instantly, e.g., within a few round-trip times (RTTs), to network failures and, consequently, bandwidth asymmetries across the topology.
@@
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SDR (Shalev et al., 2020), UE (Lu et al., 2018), Falcon (goo, 2023) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing scheme, Recycled Entropy Packet Spraying aka REPS. REPS caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
REPS does not require any specific hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). REPS requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
We extensively evaluate REPS in simulation and augment the simulation findings with real hardware results in Section 4. We deploy REPS in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, REPS consistently outperforms state-of-the-art load balancing algorithms. REPS outperforms ECMP and OPS by up to $6\times$ and $1.25\times$ in symmetric networks, and by up to $4.5\times$ and $2\times$ in asymmetric networks, outperforming OPS by as much as $100\times$ during short-term transient link failures.
2. REPS Building Blocks
In this section, we introduce the fundamental building blocks necessary to understand REPS logic. We first describe key congestion signals, such as ECN and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
@@ 2.1. Congestion Signals
In REPS, we use ECN as the congestion signal to detect path congestion due to its simplicity and widespread adoption (Hoefler et al., 2023; Nvidia, 2024). Since ECN is not marked for packets when the queue is smaller than $K_{min}$, ECN effectively filters out cases of minor queuing due to packet collisions across multiple hops, while identifying true congestion at a single bottleneck. In contrast, delay-based signals struggle to differentiate between these scenarios unless enhanced by advanced switch features, such as in-network telemetry (INT) (Wang et al., 2023).
@@
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). On current generation networks, REPS can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
@@
rely on congestion signals, such as ECN marks or packet loss, to adjust a flow’s sending rate or window with the goal of maintaining high network utilization while preventing queue build up. REPS is designed to work well with any CC algorithm as long as they support receiving and acknowledging packets out-of-order for a given message. In particular, we show later how REPS works well with EQDS, a variant of DCTCP, and a proprietary internal CC algorithm (Olteanu et al., 2022; Alizadeh et al., 2010).
@@ 2.2. Load Balancing
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). We leverage EVs in REPS to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
@@
3. REPS
Recycled Entropy Packet Spraying aka REPS is a load balancing algorithm that relies on simple and memory-efficient endpoint logic. The key idea behind REPS is straightforward: when congestion is detected on a certain path, we explore alternative paths while caching and reusing paths with little to no congestion. Specifically, REPS uses a circular buffer of a fixed size to cache EVs of uncongested paths. By design, REPS can be implemented in NIC hardware or firmware with minimal memory/area footprint. REPS does not need any switch support besides ECMP hashing and ECN marking. Placing REPS at the endpoint not only simplifies deployment but also provides a broader, end-to-end view of the network, unlike fabric-side approaches which are limited to local visibility. Algorithm 1 details pseudocode for a REPS sender receiving an ACK and detecting a failure. Algorithm 2 describes pseudocode for a REPS sender sending a data packet.
@@ 3.1. Core Logic: Path Exploration and Reuse
During the first Bandwidth-Delay Product (BDP) worth of packets of a new or idle connection, a REPS sender explores random entropies from the EVS. This exploration is necessary because, initially, there is no fresh knowledge about the network’s state. In this warm up phase, REPS operates similarly to OPS.
@@
When an ACK arrives at the sender, if it is not ECN-marked, the EV it carries is cached in the circular buffer, and its validity bit is set to 1. Otherwise, if the ACK is ECN-marked, REPS does not cache the EV and discards it. When set, the validity bit indicates that an entropy has not been used after it has been added to the circular buffer. When sending a data packet out, REPS first checks if there are any valid EVs in its buffer. In case there is any valid EV, REPS reuses the oldest valid EV from the circular buffer and resets its validity bit. Otherwise, REPS explores a random EV from the EVS.
The circular buffer in REPS ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
When a network, running without REPS, experiences any kind of transient (e.g., link flap) or persistent unrecoverable failure (e.g., a link or switch failure), it will take the system some time to recover from it: ranging from several milliseconds to update the ECMP routing group to several seconds if a reboot is needed, and more if a swap is needed (Anbiah and Sivalingam, 2022; Iselt et al., 2004).
@@
By default, REPS Core Logic would already limit utilizing failing paths since it would only recycle paths that are still active and returning ACKs without ECN. However, if the REPS buffer was empty, it would select a random EV that could potentially map to a failing path. To further improve the performance of REPS and prevent further losses, we introduce freezing mode. REPS enters this mode when detecting failures via indirect feedback from the network. REPS uses a timeout heuristic (Section 2.1) that can be enhanced with packet trimming (Appendix A) as a natural feedback from the network to detect failures along a path.
When in freezing mode, REPS:
@@
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that REPS will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we use random EVs to allow REPS to explore new paths and assess whether we detect new packets failing or not. This prevents REPS from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if REPS exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), REPS would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, REPS can be conservative and can safely enter freezing mode.
Algorithm 1 REPS logic upon ACK receive and failure detection.
@@
9:  if not repsBuffer[head].isValid then
10:   $\mathit{numberOfValidEVs}++$
@@
12:  $\mathit{repsBuffer[head].cachedEV}=\mathit{ackPacket.ev}$
13:  $\mathit{repsBuffer[head].isValid}=\textit{true}$
14:  $\mathit{head}=(\mathit{head}+1)\%\mathit{REPS\_BUFFER\_SIZE}$
@@
16:   $\mathit{isFreezingMode}=\textit{false}$
17:   $\mathit{exploreCounter}=\textit{NUM\_PKTS\_BDP}$
@@
23:   $\mathit{isFreezingMode}=\textit{true}$
24:   $\mathit{exitFreezingMode}=\textit{now()}+\textit{FREEZING\_TIMEOUT}$
@@
Algorithm 2 REPS logic on send datapath.
[…]
@@
3:  if $\mathit{numberOfValidEVs}>0$ then
4:   $\mathit{offset}=(\mathit{head}-\mathit{numberOfValidEVs})\%\mathit{REPS\_BUFFER\_SIZE}$
5:   $\mathit{repsBuffer[offset].isValid}=\textit{false}$
6:   $\mathit{numberOfValidEVs}$--
@@
8:   $\mathit{offset}=\mathit{head}$
9:   $\mathit{head}=(\mathit{head}+1)\%\mathit{REPS\_BUFFER\_SIZE}$
@@
15:  if repsBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) or exploreCounter then
16:   $dataPacket.ev=\textit{rand()}\,\%\,\mathit{EVS\_SIZE}$
17:   exploreCounter = $\max(\mathit{exploreCounter}-1,0)$
@@
19:   $dataPacket.ev=\textit{getNextEV()}$
@@
3.3. REPS Design Advantages
@@
The simplicity of REPS enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. REPS works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing.
@@
A key advantage of REPS is that it does not need to track per-EV metrics and statistics. As will be discussed in Section 4.5.2, achieving good performance with OPS requires a relatively large EVS. If OPS were to maintain metrics for each EV, the memory overhead would be excessive for a hardware NIC implementation, e.g., 64 Kib to store 1 bit per entropy value for an EVS with 64K EVs. However, REPS only needs a fixed number of bytes in memory. More specifically, as detailed in Table 1, REPS requires only around $\approx 25$ bytes. Moreover, even when constrained to a small EVS, REPS is still able to perform well (Section 4.5.2), which can further reduce REPS’ memory footprint by 1 byte since Table 1 assumed 16 bits per EV.
There is a subtle observation as to why REPS can achieve a great performance without needing a lot of state: while the REPS buffer is used to cache good entropies, it is really only useful in certain scenarios like when receiving a burst of ACKs or during freezing mode. In reality, most of REPS’ state is on the wire, stored in the inflight data and ACK packets, which will inform REPS about the good paths in the network.
@@
Table 1. Per-connection memory footprint of REPS.
@@
The general approach of REPS is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.